OpenStudy (stephanieelizzz):

The height in feet, h, reached by a projectile after t seconds is given by the function: h(t) = -16t2 + 110t + 5 During what time interval did the projectile reach an altitude of at least 135 ft?

OpenStudy (mathstudent55):

You want the height to be at least 135 ft. Is "at least 135 ft" greater than of equal to 135 ft or is it less than or equal to 135 ft?

OpenStudy (stephanieelizzz):

It would be greater than or equal to 135?

OpenStudy (stephanieelizzz):

@mathstudent55

OpenStudy (mathstudent55):

Correct. Now set your expression greater than or equal to 135 and solve for t.

OpenStudy (mathstudent55):

\(-16t^2 + 110t + 5 \ge 135\)

OpenStudy (stephanieelizzz):

Would you factor out first and then put them equal to zero?

OpenStudy (mathstudent55):

Subtract 135 from both sides. Then treat it as an equation by setting it equal to zero. Use the quadratic formula to find values of t.

OpenStudy (mathstudent55):

\(-16t^2 + 110t - 130 \ge 0\) Solve for t: \(-16t^2 + 110t - 130 = 0\)

OpenStudy (mathstudent55):

You can divide both sides by -2, if you'd like to work with smaller numbers. Then use the quadratic formula. \(8t^2 - 55t + 65 = 0\)

OpenStudy (mathstudent55):

\(\Large x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \) \(\Large t = \dfrac{-(-55) \pm \sqrt{(-55)^2 - 4(8)(65)}}{2(8)} \) \(\Large t = \dfrac{55 \pm \sqrt{3025 - 2080}}{16} \) \(\Large t = \dfrac{55 \pm \sqrt{945}}{16} \) \(\Large t = 1.52\) or \(t = 5.36\)

OpenStudy (mathstudent55):

The interval is between the two times above.